Please do questions 13-15 and show all work please!

13) A magazine editor must choose 4 short stories for this month’s issue from 30 submissions. In how many ways can the editor choose this month’s stories?

14) When five cards are drawn without replacement from a standard deck of 52 cards, find the probability of obtaining exactly two kings.

15) Find the probability of selecting six winning numbers from 1 to 51.
Answer #1

13) number of ways to choose 4 stories from 30 =30C4 =30!/(26!*4!)=27405

14)number of ways to draw 5 cards =52C5 =2598960

number of ways to draw 2 kings from 4 cards and remaining 3 cards from rest of 48 cards

=4C2*48C3 =6*17296=103776

therefore probability=103776/2598960

15)(for this we have not given information about the game.. please revert)

probability of selecting six winning numbers from 1 to 51 =1/51C6 =1/18009460